# Metron: Holistic Performance Evaluation Framework for LLM Inference Systems

### v1

@@ Abstract
Serving large language models (LLMs) in production can incur substantial costs, which has prompted recent advances in inference system optimizations. Today, these systems are evaluated against conventional latency and throughput metrics (eg. TTFT, TBT, Normalised Latency and TPOT). However, these metrics fail to fully capture the nuances of LLM inference, leading to an incomplete assessment of user-facing performance crucial for real-time applications such as chat and translation. In this paper, we first identify the pitfalls of current performance metrics in evaluating LLM inference systems. We then propose Metron, a comprehensive performance evaluation framework that includes fluidity-index– a novel metric designed to reflect the intricacies of the LLM inference process and its impact on real-time user experience. Finally, we evaluate various existing open-source platforms and model-as-a-service offerings using Metron, discussing their strengths and weaknesses. Metron is available at github.com/project-metron/metron.
@@ 1 Introduction
To address the limitations of existing metrics, we introduce Metron, a comprehensive framework for evaluating user-facing performance in LLM inference. At its core are two novel metrics: fluidity-index and fluid token generation rate, designed to capture the nuances of real-time, streaming LLM interactions. We apply this framework to conduct an extensive performance evaluation of both open-source and proprietary LLM inference systems, revealing their strengths and weaknesses.
@@
We have open-sourced Metron at github.com/project-metron/metron, aiming to establish a standard for user-centric performance evaluation in the rapidly evolving landscape of LLM inference systems and frameworks.
@@ 3.2 Desirable Properties of Evaluation Framework
4 Metron: Design and Implementation
@@
Figure 4: When a request arrives in the system, Metron sets the deadlines for all future tokens. If a token is produced before the set deadline, the slack is carried forward and serves as a buffer for future tokens. When a token arrives late, the system gets penalized for all the missed deadlines, and the subsequent deadlines are reset to account for autoregressive decoding process.
@@
Picking the larget latencies. Two important constants in the fluidity-index metric is the target latencies for the first generated token (prefill) and subsequent (decode) tokens denoted by $D_{p}$ and $D_{d}$ above. As discussed in §3.1, the prefill time increases quadratically in the size of input prompt. Therefore, setting a static value for $D_{p}$ is impractical as the prompt length varies significantly in production scenarios. LMSys-Chat-1M  [31] has 417 median number of prefill tokens and 1418 prefill tokens at $90^{th}$ percentile. Therefore, in Metron, we propose $D_{p}$ to be a function of prompt length. We use a reasonable open-source system like vLLM [14] as baseline to benchmark the prefill time as a function of input length by profiling the request in isolation, in the absence of any other variables like scheduling delays. We repeat this for 10 requests (tunable) and fit a curve through the observed points to obtain a $D_{p}$ target as a function of input length. Note that, this is a recommendation on how to set the prefill latency target; different systems may observe different performance trends due to their implementation, optimizations, and kernels used. Therefore, the process of identifying this prefill latency curve should be repeated on the system being evaluated to set practical latency targets.
Picking the decode latency target $D_{d}$ is fairly intuitive and straightforward. It depends on the application being evaluated. We define three targets in Metron– a strict target for interactive applications like chat (25ms), medium target for medium-priority users (50ms) and a relaxed target for low-priority users (100ms). These numbers in Metron can be picked based on production needs.
@@ 4.2 Evaluation workflow and implementation
Metron, provides a standardized evaluation workflow for both proprietary and open-source LLM inference frameworks. Metron provides two evaluation recipes as described below:
Black-box Evaluation. For an LLM inference API endpoint, Metron performs black-box evaluation by hitting the server with a set of requests with diverse prompt lengths, and tracks checkpoints such as the timestamps when each output token got generated, which allows it to calculate several metrics such as TTFT, TBT, and TPOT. Moreover, given a target threshold value for fluidity-index metric, such as – 99% of the requests have their fluidity-index of at-least 0.9, Metron infers the minimum TBT deadline that can satisfy this constraint. This allows us to obtain the maximum stable token throughput (fluid token generation rate) that can be served to the user in glitch-free manner.
Capacity Evaluation. Typically, while deploying an LLM inference service, the operator needs to determine the minimum number of GPUs required to serve the expected userbase with predefined service quality requirements. To aid this process, Metron provides a capacity evaluation module, which evaluates a system with different request loads to identify the maximum capacity each replica can provide while meeting the fluidity-index SLO requirements.
@@ 5 Evaluation
In this section we demonstrate the effectiveness of Metron in holistically evaluating the performance of different LLM inference systems both open source frameworks and proprietary offerings.
@@ 5.1 Evaluating Public APIs
In this section, we demonstrate the effectiveness of Metron to benchmark public API endpoints. Metron performs black-box analysis on these systems to characterize their performance under various configurations. We evaluate three proprietary systems with API-only access: Anyscale [1], Groq [5], and Fireworks [4], across two models – a dense model LLaMA3-70B[9], and a MoE model Mixtral-8x7B [24]. We use a custom workload with varying prefill length (between 256 and 8k and maximum tokens to generate set to 256. Since the performance of public APIs can change throughout the day depending on request traffic and other factors, we run Metron once every hour for 24 hours to accommodate varying nature of traffic throughout the day.
@@ 5.2 Evaluating Open Source Systems
We now demonstrate the effectiveness of Metron in setting SLOs for deployment operators and capacity planning. We evaluate vLLM [14] and Sarathi-Serve [17] (via vLLM with chunked-prefill feature turned on). on LLaMA3-8B [9, 29] on a H100. We use rope-scaling to support prompt token lengths longer than 8192. Requests are randomly sampled from the Arxiv-Summarization [18] dataset that represents long context workloads.
@@
Results. Earlier in  §3.1 we compared the vLLM and Sarathi-Serve at a high load, where we observed that the tail TBT-based throughput for vLLM is 3$\times$ worse than Sarathi due to huge generation stalls, while the TPOT based throughput shows these systems at par. fluidity-index shows the true difference in throughput between these systems. Next, we compare these systems under a strict TBT SLO. We use fluidity-index to define the service SLO as – 99% of requests should have less than 10% deadline miss rate with 25ms target TBT. Metron then finds the maximum request load (QPS) at which the SLOs can be maintained. We also consider TPOT and P99 TBT based SLOs as baselines with the target latency of 25ms. Here, Sarathi has a 2x lower token throughput compared to vLLM as measured by tail TBT-based metric, in contrast to Figure 3. At a budget of 25ms, almost all the mixed batches created by Sarathi-Serve with prefill chunks violate the latency SLO threshold.  Figure 6b, shows the distribution of deadline miss rate when both the systems are operating at the capacity. Also, note that in this setting, the fluidity-index based capacity is the same at 0.6 QPS. We find that vLLM has a higher deadline miss rate at lower percentile. This is expected because, vLLM will have to ingest prefills in their entirety from time to time and whenever it does so, the TBT deadline is breached. In contrast, with chunked prefills Sarathi-Serve achieves fewer deadline misses and higher fluidity.
@@ 6 Discussion
Metron provides an evaluation framework for LLM inference using fluidity-index metric that tracks the missed deadlines per request. We now discuss the challenges that we leave to future work.
@@
Next, we observe that we need to provide a small scheduling slack in deadline computations as discussed in §4.1. We pick this value based on our empirical observations; we leave it to future work to systematically set a scheduling slack. Finally, open-source systems have various performance tuning knobs; for e.g., chunk size in Sarathi-Serve, block size in vLLM etc. Metron currently does not explore or auto-tune such parameters; it expects the users to set the configuration parameters while evaluating across two systems.
@@ 8 Conclusion
Evaluating LLM inference systems is a challenging problem due to the unique characteristics of autoregressive decode process. We presented a detailed analysis of existing evaluation metrics and their pitfalls. To address their shortcomings, we introduce Metron– a holistic LLM evaluation framework that instantiates a novel approach comprised of a novel fluidity-index based approach to evaluating LLM inference systems in a user-facing manner. We then show how Metron can be leveraged to evaluate both open-source and proprietary model serving systems. Metron is aimed to serve as a standard evaluation suite for LLM inference systems.

### v2

# Etalon: Holistic Performance Evaluation Framework for LLM Inference Systems

@@ Abstract
Serving large language models (LLMs) in production can incur substantial costs, which has prompted recent advances in inference system optimizations. Today, these systems are evaluated against conventional latency and throughput metrics (eg. TTFT, TBT, Normalised Latency and TPOT). However, these metrics fail to fully capture the nuances of LLM inference, leading to an incomplete assessment of user-facing performance crucial for real-time applications such as chat and translation. In this paper, we first identify the pitfalls of current performance metrics in evaluating LLM inference systems. We then propose Etalon, a comprehensive performance evaluation framework that includes fluidity-index– a novel metric designed to reflect the intricacies of the LLM inference process and its impact on real-time user experience. Finally, we evaluate various existing open-source platforms and model-as-a-service offerings using Etalon, discussing their strengths and weaknesses. Etalon is available at github.com/project-etalon/etalon.
@@ 1 Introduction
To address the limitations of existing metrics, we introduce Etalon, a comprehensive framework for evaluating user-facing performance in LLM inference. At its core are two novel metrics: fluidity-index and fluid token generation rate, designed to capture the nuances of real-time, streaming LLM interactions. We apply this framework to conduct an extensive performance evaluation of both open-source and proprietary LLM inference systems, revealing their strengths and weaknesses.
@@
We have open-sourced Etalon at github.com/project-etalon/etalon, aiming to establish a standard for user-centric performance evaluation in the rapidly evolving landscape of LLM inference systems and frameworks.
@@ 3.2 Desirable Properties of Evaluation Framework
4 Etalon: Design and Implementation
@@
Figure 4: When a request arrives in the system, Etalon sets the deadlines for all future tokens. If a token is produced before the set deadline, the slack is carried forward and serves as a buffer for future tokens. When a token arrives late, the system gets penalized for all the missed deadlines, and the subsequent deadlines are reset to account for autoregressive decoding process.
@@
Picking the larget latencies. Two important constants in the fluidity-index metric is the target latencies for the first generated token (prefill) and subsequent (decode) tokens denoted by $D_{p}$ and $D_{d}$ above. As discussed in §3.1, the prefill time increases quadratically in the size of input prompt. Therefore, setting a static value for $D_{p}$ is impractical as the prompt length varies significantly in production scenarios. LMSys-Chat-1M  [31] has 417 median number of prefill tokens and 1418 prefill tokens at $90^{th}$ percentile. Therefore, in Etalon, we propose $D_{p}$ to be a function of prompt length. We use a reasonable open-source system like vLLM [14] as baseline to benchmark the prefill time as a function of input length by profiling the request in isolation, in the absence of any other variables like scheduling delays. We repeat this for 10 requests (tunable) and fit a curve through the observed points to obtain a $D_{p}$ target as a function of input length. Note that, this is a recommendation on how to set the prefill latency target; different systems may observe different performance trends due to their implementation, optimizations, and kernels used. Therefore, the process of identifying this prefill latency curve should be repeated on the system being evaluated to set practical latency targets.
Picking the decode latency target $D_{d}$ is fairly intuitive and straightforward. It depends on the application being evaluated. We define three targets in Etalon– a strict target for interactive applications like chat (25ms), medium target for medium-priority users (50ms) and a relaxed target for low-priority users (100ms). These numbers in Etalon can be picked based on production needs.
@@ 4.2 Evaluation workflow and implementation
Etalon, provides a standardized evaluation workflow for both proprietary and open-source LLM inference frameworks. Etalon provides two evaluation recipes as described below:
Black-box Evaluation. For an LLM inference API endpoint, Etalon performs black-box evaluation by hitting the server with a set of requests with diverse prompt lengths, and tracks checkpoints such as the timestamps when each output token got generated, which allows it to calculate several metrics such as TTFT, TBT, and TPOT. Moreover, given a target threshold value for fluidity-index metric, such as – 99% of the requests have their fluidity-index of at-least 0.9, Etalon infers the minimum TBT deadline that can satisfy this constraint. This allows us to obtain the maximum stable token throughput (fluid token generation rate) that can be served to the user in glitch-free manner.
Capacity Evaluation. Typically, while deploying an LLM inference service, the operator needs to determine the minimum number of GPUs required to serve the expected userbase with predefined service quality requirements. To aid this process, Etalon provides a capacity evaluation module, which evaluates a system with different request loads to identify the maximum capacity each replica can provide while meeting the fluidity-index SLO requirements.
@@ 5 Evaluation
In this section we demonstrate the effectiveness of Etalon in holistically evaluating the performance of different LLM inference systems both open source frameworks and proprietary offerings.
@@ 5.1 Evaluating Public APIs
In this section, we demonstrate the effectiveness of Etalon to benchmark public API endpoints. Etalon performs black-box analysis on these systems to characterize their performance under various configurations. We evaluate three proprietary systems with API-only access: Anyscale [1], Groq [5], and Fireworks [4], across two models – a dense model LLaMA3-70B[9], and a MoE model Mixtral-8x7B [24]. We use a custom workload with varying prefill length (between 256 and 8k and maximum tokens to generate set to 256. Since the performance of public APIs can change throughout the day depending on request traffic and other factors, we run Etalon once every hour for 24 hours to accommodate varying nature of traffic throughout the day.
@@ 5.2 Evaluating Open Source Systems
We now demonstrate the effectiveness of Etalon in setting SLOs for deployment operators and capacity planning. We evaluate vLLM [14] and Sarathi-Serve [17] (via vLLM with chunked-prefill feature turned on). on LLaMA3-8B [9, 29] on a H100. We use rope-scaling to support prompt token lengths longer than 8192. Requests are randomly sampled from the Arxiv-Summarization [18] dataset that represents long context workloads.
@@
Results. Earlier in  §3.1 we compared the vLLM and Sarathi-Serve at a high load, where we observed that the tail TBT-based throughput for vLLM is 3$\times$ worse than Sarathi due to huge generation stalls, while the TPOT based throughput shows these systems at par. fluidity-index shows the true difference in throughput between these systems. Next, we compare these systems under a strict TBT SLO. We use fluidity-index to define the service SLO as – 99% of requests should have less than 10% deadline miss rate with 25ms target TBT. Etalon then finds the maximum request load (QPS) at which the SLOs can be maintained. We also consider TPOT and P99 TBT based SLOs as baselines with the target latency of 25ms. Here, Sarathi has a 2x lower token throughput compared to vLLM as measured by tail TBT-based metric, in contrast to Figure 3. At a budget of 25ms, almost all the mixed batches created by Sarathi-Serve with prefill chunks violate the latency SLO threshold.  Figure 6b, shows the distribution of deadline miss rate when both the systems are operating at the capacity. Also, note that in this setting, the fluidity-index based capacity is the same at 0.6 QPS. We find that vLLM has a higher deadline miss rate at lower percentile. This is expected because, vLLM will have to ingest prefills in their entirety from time to time and whenever it does so, the TBT deadline is breached. In contrast, with chunked prefills Sarathi-Serve achieves fewer deadline misses and higher fluidity.
@@ 6 Discussion
Etalon provides an evaluation framework for LLM inference using fluidity-index metric that tracks the missed deadlines per request. We now discuss the challenges that we leave to future work.
@@
Next, we observe that we need to provide a small scheduling slack in deadline computations as discussed in §4.1. We pick this value based on our empirical observations; we leave it to future work to systematically set a scheduling slack. Finally, open-source systems have various performance tuning knobs; for e.g., chunk size in Sarathi-Serve, block size in vLLM etc. Etalon currently does not explore or auto-tune such parameters; it expects the users to set the configuration parameters while evaluating across two systems.
@@ 8 Conclusion
Evaluating LLM inference systems is a challenging problem due to the unique characteristics of autoregressive decode process. We presented a detailed analysis of existing evaluation metrics and their pitfalls. To address their shortcomings, we introduce Etalon– a holistic LLM evaluation framework that instantiates a novel approach comprised of a novel fluidity-index based approach to evaluating LLM inference systems in a user-facing manner. We then show how Etalon can be leveraged to evaluate both open-source and proprietary model serving systems. Etalon is aimed to serve as a standard evaluation suite for LLM inference systems.